please help

Three forces act simultaneously on a 2.70-kg block. For clarity the drawings show the forces separately. The block starts from rest and, as a result of the combinedaction of the forces, undergoes a displacement s, which has a magnitude of 6.50 m. What is the final speed of the block?
Answer #1
No drawing given.
answered by: Lusi
Answer #2
v^2=2as
answered by: malik jackson
Answer #3

displacement = 6.5 m
speed =√2as

a =F/m

u have not given figure u find out net force from figure and then find accelration

substitute in above formula u ll get answer

answered by: Benjua
Answer #4
i need the drawings !!
answered by: importance of writing a new mission statement after taking o
Answer #5
find the resultant force from drawing
now use formula
a =F/m
to get acceleration then v =u +at
gives the final velocity
answered by: ericca porch